Determine whether u and v are orthogonal vectors, make an acute or obtuse angle:(1.1) u =<1,3,−2 >, v =<−5,3,2 >.

(1.2) .u =<1,−2,4 >, v =<5,3,7 >.


1
Expert's answer
2021-06-25T05:29:11-0400

1.1

"u=(1,3,-2), v=(-5,3,2)\\
uv=-5+9-4=0\\"

u orthogonal v


1.2

"u=(1,-2,4), v=(5,3,7)\\
uv=5-6+28=27>0"

u is not orthogonal v

"||u||=\sqrt{1+4+16}=\sqrt{21}\\
||v||=\sqrt{25+9+49}=\sqrt{83}\\
\cos\phi=\frac{uv}{||u||\cdot||v||}=\frac{27}{\sqrt{21}\cdot\sqrt{83}}>0"

acute angle